Calculate the wavelength of thermal neutrons at 20 degree Celsius





1
Expert's answer
2021-11-07T19:26:57-0500

Begin with the energy:


"E=\frac 32 kT,\\\space\\
\lambda=\frac h{\sqrt{2mE}}=\frac h{\sqrt{3mkT}}=147\text{ pm}."